OpenStudy (anonymous):

A simple problem., if abc = 10 acd = 5 abd = 45 bcd = 12 what is a + c?

OpenStudy (anonymous):

i'm useless here .. brb

OpenStudy (anonymous):

No offense but: don't be expecting an answer on OS if you have no dedication of trying in the first place.

OpenStudy (callisto):

@Noliec I think shai is not that kind of person

OpenStudy (anonymous):

If that is the case then I apoligize.

OpenStudy (apoorvk):

@Noliec you are probably mistaken. She posts some really good problems, and is way more genius than a lot of us (atleast me) :)

OpenStudy (anonymous):

This could be relatively easy solved with matrices.

OpenStudy (anonymous):

multiply them all together to get (abcd)^3. from there you can figure out what a, b, c, and d are individually.

OpenStudy (anonymous):

@joemath314159 that can help a lot .. thanks

OpenStudy (anonymous):

then abcd = 30 .. ?

OpenStudy (callisto):

I think so

OpenStudy (anonymous):

Yep, 30

OpenStudy (anonymous):

abc = 10, abcd/d = 30/d = 10 --> 30 = 10 d --> d = 3

OpenStudy (anonymous):

@Noliec i want to learn, that's why i'm here in OS

OpenStudy (anonymous):

I am sorry, I misinterpreted your first reply to this question. :)

OpenStudy (anonymous):

in this case, there's no need to solve for the values of b and d right?

OpenStudy (anonymous):

i think i figure it out now,.. hmmm

OpenStudy (callisto):

You can do so, but not necessary, I think...

OpenStudy (anonymous):

So a = 5/2, b = 6, c = 2/3, d = 6. a+c=19/6

OpenStudy (anonymous):

@Recursing you're smart! Next problem .. lol